# ColorVideoVDP: A visual difference predictor for image, video and display distortions

### v1

@@ 3.7 Pooling
$\left\|f(v)\right\|_{p,v}=\left(\sum_{v}|f(v)|^{p}\right)^{\nicefrac{{1}}{{p}}}\,.$ (15)
$w_{c}$ is the weight associated with each channel. The main advantage of pooling differences first across pixels is that we avoid the expensive step of the synthesis of the distortion map from the multiple levels of the Laplacian pyramid. It should be noted that the differences across pixels are normalized by the number of pixels in each band, $N_{b}$ — we do not want the bands represented with more pixels in the pyramid to contribute more to image quality. The pooling exponents $\beta_{x}$ and $\beta_{b}$ were optimized in FovVideoVDP, but we found that such an optimization is unstable (because of exploding gradients) and unnecessary. Instead, we set $\beta_{x}=2$ and $\beta_{f}=2$ to represent the energy summation across the spatial dimensions and frames Watson et al. [1983]. $\beta_{b}$ and $\beta_{c}$ are set to 4, which is representative of the summation across channels Quick et al. [1978]. Using exponents greater than 1 roughly corresponds to a ”winner-take-all” strategy, in which the strongest visual differences have the highest impact on the pooled value.
3.8 JOD regression
Figure 10: Interpretation of the JOD scores. (a) JOD quality scores are normally distributed random variables — the reported JOD value is the mean of the distribution, which models that a certain portion of the population will rate the quality as higher or lower than the reported mean value. (b) The standard deviation of the distribution is selected in such a way that when the difference between the means is 1 JOD ($q_{2}-q_{1}=1$), 75% of the population will select the condition with the higher JOD value as better. This allows us to interpret quality differences expressed in JOD units.
@@ 3.8 JOD regression
$Q_{\textrm{JOD}}=10-\alpha_{\textrm{JOD}}\,(D_{\textrm{pooled}})^{\beta_{\textrm{JOD}}}\,,$ (16)
@@ 3.9 Image quality
$D_{\textrm{pooled}}^{\textrm{image}}=k_{\text{I}}\left\|w_{c}\left\|\frac{1}{N_{b}^{\nicefrac{{1}}{{\beta_{x}}}}}\left\|\hat{D}_{b,c,f}(\bm{x})\right\|_{\beta_{x},\bm{x}}\right\|_{\beta_{b},b}\right\|_{\beta_{c},c}\,.$ (17)
@@
3.10 Visualization — heatmaps and distograms
Figure 11: Examples of distogram visualization of artifacts. The values on the y-axis denote spatial frequency in cpd. The higher (more yellowish) values denote stronger distortion. The low-frequency waveguide nonuniformity (left) is shown to be visible mostly in the base band (BB) and low frequency bands of both achromatic (A-sust) and chromatic (RG, YV) channels. The artifacts shows also as low-frequency flicker in the achromatic transient channel (A-trans). The (per-pixel) light source nonuniformity causes high-frequency shimmer, which is visible as high frequency distortions in the distogram on the right.
@@ 3.10 Visualization — heatmaps and distograms
$H_{b}(\bm{x})=\left\|w_{c}\hat{D}_{b,c,f}(\bm{x})\right\|_{\beta_{c},c}$ (18)
@@
Another visualization, which we name a distogram, lets us present video distortions within each channel and spatial frequency band, all in a single diagram. Two examples for two distortions from the XR-DAVID dataset are shown in Fig. 11. The examples show that depending on the characteristics of the distortion, the artifacts will show in either low- or high-frequency bands, in achromatic or chromatic channels. This visualization lets us interpret and explain the single-valued JOD scores assigned to the distorted videos.
3.11 Implementation details
@@
4 XR Display Artifact Video Dataset (XR-DAVID)
@@
To address this gap in quantitative data for XR distortions, we enumerated the most relevant artifacts for our study. We placed special emphasis on distortions that have a color or temporal component, as these are not well served by existing metrics and are the focus of ColorVideoVDP . We used this dataset of distortions to conduct a large-scale subjective study, collecting JOD quality scores which are suitable for calibration of ColorVideoVDP .
Figure 12: Physical hardware setup for the XR-DAVID data collection. Participants used a chinrest to fix their distance from the display at 0.73 m (two display heights) and used a 3-key keyboard to record their choices.
4.1 Experimental setup
@@
[…]
@@
We conducted the study using paid external participants. After two pilots (N=5 each), 77 naïve users took part in 1 hour long sessions of our main study, which included training, data gathering, and a short break half-way through. All participants signed informed consent forms, and the study was approved by an external ethics committee. Participants were screened via an Ishihara color test Hardy et al. [1945] to ensure normal color vision.
@@
Participants performed a side-by-side pairwise comparison task answering which of the two versions of a video was less distorted. This method was preferred to alternatives (e.g., direct ratings on the mean-opinion-score scale) as pairwise comparisons have been shown to produce more accurate results Zerman et al. [2018] by simplifying the task in each trial. Active sampling using the ASAP method Mikhailiuk et al. [2021] was used to optimize information gain from each comparison, making it possible to explore a large number of artifacts without prohibitively increasing the number of pairwise comparisons. We allowed comparisons across different distortion types and levels but not across different video contents. A training session preceded the main experiment. In that session, the participants compared each reference video to its distorted version. A different distortion type was used in each trial. The training session was meant to familiarize the participants with the reference content and distortions. The results of the main experiment were scaled to a unified perceptual scale in just-objectionable-difference units (JODs) using the pwcmp software suite Perez-Ortiz et al. [2019].
4.2 Distortions
@@
Figure 13: MTF Curves corresponding to the Blur artifact at level 1, 2, and 3.
@@
AR displays with diffractive waveguides such as Microsoft Hololens, Magic Leap, and Snap Spectacles Ooi and Dingliana [2022] exhibit a characteristic spatially-varying color nonuniformity. This nonuniformity is typically low-frequency (less than  1 cycle per degree). In addition, it is heavily dependent on pupil position within the eye box, which can vary depending on the user’s fixation and eye movements. In particular, if a user keeps their gaze fixed on an object in the AR content while moving their head (engaged in what is typically termed a vestibulo-ocular reflex or VOR movement), they may notice its color shifting as the position of the eye changes in relation to the waveguide. In order to simulate the artifact, first, two nonuniformity patterns were obtained from empirical waveguide transmission data. Second, base variation maps were generated for the two patterns by cleaning noise in data and normalizing variations across channels. The variations were also scaled such that they appear to have similar magnitude across the two patterns. Finally, the amplitude of variations in the base map was varied to produce different artifact levels. The multipliers of the amplitude variations were 0.15, 0.3, and 0.4 for artifact levels 1, 2, and 3 respectively. To simulate changes with eye position, the nonuniformity pattern transitions from the first base pattern to the second and back over the course of the video. This transition is repeated 4 times for level 2, and 7 times for level 3. Fig. 14 illustrates this temporally varying waveguide nonuniformity artifact.
Figure 14: Examples of waveguide nonuniformity level 3 at different frames, using Business scene (cropped for better visibility). nonuniformity changes from one pattern (left) to another (right).
@@
In AR displays suffering from waveguide nonuniformity across different exit pupil positions, as described above, a dynamic correction algorithm can be implemented to invert the distortion, reducing spatial nonuniformity. This algorithm would be dependent on gaze location, thus relying on estimated positions from eye trackers, which may be inaccurate. Namely, inaccurate eye tracker readings will cause a wrong estimate of the pupil position and temporal color artifacts (by either not correcting perfectly for distortions or performing an imprecise correction for a distortion pattern that is not present). To simulate this, the precision error was randomly generated per frame from a Gaussian distribution with a standard deviation of 0.27°, 0.39°, and 0.58° for artifact levels 1, 2, and 3 respectively. Simulated precision errors are in alignment with existing VR headsets currently in the market Schuetz and Fiehler [2022].
@@ Results
Figure 15: The results of our display artifact study aggregated over all users and scenes are shown. Quality scores are expressed in JODs, counting down from 10 by convention. Values above the nominal reference are the result of chance selection for artifacts that appear very subtle. Error bars represent 95% confidence intervals, obtained via bootstrapping.
5 Metric validation
Table 2: Datasets used for training and testing.
@@ 5 Metric validation
UPIQ Mikhailiuk et al. [2022] Train & test SDR/HDR images 84 4159 34 distortion types
LIVE HDR Shang et al. [2022] Test HDR video 21 210 H.265, bicubic upscaling
LIVE VQA Seshadrinathan et al. [2010] Test HDR video 10 150 H.264, MPEG-2, transmission
KADID-10k Lin et al. [2019] Test SDR images 81 10125 25 distortion types
@@
We split the datasets into those used solely for testing and those used for both training and testing. The five selected datasets are listed in Table 2. The two datasets used for both training and testing are XR-DAVID, explained in detail in Sec. 4, and UPIQ Mikhailiuk et al. [2022]. We selected the UPIQ dataset because it contains a large collection of both SDR and HDR images (over 4000) and is scaled in JOD units, similar to XR-DAVID. To test our metric on unseen datasets (cross-dataset validation), we chose LIVE HDR because it is a modern dataset representative of video streaming applications, KADID-10k because of its size (over 10k images), and LIVE VQA because it is widely used to test video quality metrics.
@@
Training a video quality metric is problematic as a very large amount of data is used to predict a single quality value. For example, 60 frames of 4K video requires the processing of almost 500 million pixels to infer just a single quality score. This makes gradient computation (through backpropagation) infeasible because of the memory requirements. Previous work dealt with this problem in several ways: some metrics were designed to extract low dimensional features from a video and then train a regression mapping those features to quality scores Li et al. [2016b]. This approach, however, does not allow training or tuning the feature extraction stage. Another group of methods operated on patches of limited resolution (e.g. 64$\times$64), assuming that the quality is the same across the entire image Prashnani et al. [2018]. This assumption, however, can be easily proved wrong for localized distortions or for content in which the effect of contrast masking varies across an image. Other metrics used numerical gradient computation Mantiuk et al. [2021], which, however, becomes too expensive when optimizing a large number of parameters.
@@
For the end-to-end training, we computed the full analytical gradient of the remaining parameters using two techniques. Firstly, we utilized gradient checkpointing Chen et al. [2016] in PyTorch, by rerunning a forward pass for each checkpointed segment during backward propagation. This technique trades some speed for reduced memory requirements. We inserted a checkpoint after each block of frames, where the block size was determined based on the available GPU memory. Secondly, during training, we randomly sampled 0.5-second-long sequences from each video clip (full sequences were used for testing). This approach improved training time and introduced a form of data augmentation. The feature-based and end-to-end training were run in an interleaved manner, with 50 epochs of (fast) feature-based training followed by a single epoch of end-to-end training. This lets us jointly train all the parameters of the metric. The trained parameters of the metric can be found in Table 3.
Table 3: The trained parameters of the ColorVideoVDP . The parameters of the cross-channel masking can be found in Fig. 9.
Model component
Parameters
Contrast sensitivity
$S_{\text{corr}}=0.9683$
Masking
$p=2.264$, $q_{c}=[1.303,2.889,3.681,3.589]$, $k_{\text{C}}=366.6$, $\sigma_{\text{sp}}=3$,
@@
Pooling
$k_{\text{I}}=0.5779$, $w_{c}=[1,1,1,0.8081]$
JOD regression
$\alpha_{\text{JOD}}=0.04396$, $\beta_{\text{JOD}}=0.9302$
5.1 Comparison with other metrics
@@
As our datasets include both SDR and HDR content, we need to ensure that those are handled correctly by all the metrics. For metrics that do not work with colorimetric data and instead operate on display-encoded (SDR) pixel values, we ran the evaluations using the original SDR pixel values. In the case of HDR content, we employed the perceptually-uniform transform (PU21) Mantiuk and Azimi [2021] to encode the pixel values. The metrics that operate on colorimetric data (FovVideoVDP, HDR-VDP-3) on the other hand, were supplied the absolute luminance values, computed by the display model (Sec. 3.1).
@@
The results, shown in Fig. 16, indicate a substantial gain in performance of ColorVideoVDP  over the second-best metric — VMAF. Image metrics that consider color — FSIMc and VSI — performed better than expected on XR-DAVID, but rather poorly on LIVEHDR video dataset. The color difference formulas and their spatial extensions performed worse than image and quality metrics. The non-reference metrics — NIQE, PIQE, and BRISQUE — show a very small correlation with subjective judgments. Other performance indices (SROCC and PLCC) and detailed results can be found in the HTML reports included in the supplementary. Overall, while it is possible to find a metric that performs well on a selected dataset, such as STRRED on LIVEVQA, ColorVideoVDP  offers good performance across all datasets with a wide variety of distortions and content. This could be explained by no other metric offering the same set of capabilities as ColorVideoVDP ; only our metric models spatiotemporal color vision and accounts for the display model.
Figure 16: ColorVideoVDP  compared with existing metrics, listed in Table 1. The prediction errors are reported for testing datasets (LIVEVQA, LIVEHDR and KADID) and the test portion of UPIQ and XR-DAVID. The error bars denote 95% confidence intervals. Other performance metrics (SROCC and PLCC) and detailed reports can be found in the supplementary HTML report.
5.2 Ablations
Figure 17: Ablation studies on the contrast encoding and masking models used in ColorVideoVDP . The notation is the same as in Fig. 16. The masking models used in the ablations are explained in the supplementary.
We grouped ablation studies into those used to determine suitable contrast encoding and masking models, and into those used to test the importance of each component of the metric. The metric parameters were refitted for each ablation, as explained above. First, we tested a combination of two contrast encodings and three masking functions, all explained in detail in the supplementary document. The results of those ablations, shown in Fig. 17, clearly indicate that the mutual masking model with multiplicative contrast encoding offers much better performance than the alternatives. The results indicate that the multiplicative contrast encoding is necessary for mutual masking and transducer models. As mentioned in Sec. 3.6, such encoding can unify results across different spatial frequencies Daly [1993] and color directions Cass et al. [2009]; Switkes et al. [1988]. Although the contrast transducer is one of the best-established masking models Watson and Solomon [1997], which accounts for facilitation and performs well on selected datasets Alam et al. [2014], it did not perform well as a part of the quality metric. The similarity formula exhibits masking properties and is used in many metrics, such as (MS-)SSIM Wang et al. [2003], but it did not result in acceptable performance when integrated into our metric.

### v2

@@ 3.7. Pooling
(15) $\left\|f(v)\right\|_{p,v}=\left(\sum_{v}|f(v)|^{p}\right)^{\nicefrac{{1}}{{p}}}\,.$
$w_{c}$ is the weight associated with each channel. The main advantage of pooling differences first across pixels is that we avoid the expensive step of the synthesis of the distortion map from the multiple levels of the Laplacian pyramid. It should be noted that the differences across pixels are normalized by the number of pixels in each band, $N_{b}$ — we do not want the bands represented with more pixels in the pyramid to contribute more to image quality. The pooling exponents $\beta_{x}$ and $\beta_{b}$ were optimized in FovVideoVDP, but we found that such an optimization is unstable (because of exploding gradients) and unnecessary. Instead, we set $\beta_{x}=2$ and $\beta_{f}=2$ to represent the energy summation across the spatial dimensions and frames (Watson et al., 1983). $\beta_{b}$ and $\beta_{c}$ are set to 4, which is representative of the summation across channels (Quick et al., 1978). Using exponents greater than 1 roughly corresponds to a ”winner-take-all” strategy, in which the strongest visual differences have the highest impact on the pooled value.
3.8. JOD regression
Figure 10. Interpretation of the JOD scores. (a) JOD quality scores are normally distributed random variables — the reported JOD value is the mean of the distribution, which models that a certain portion of the population will rate the quality as higher or lower than the reported mean value. (b) The standard deviation of the distribution is selected in such a way that when the difference between the means is 1 JOD ($q_{2}-q_{1}=1$), 75% of the population will select the condition with the higher JOD value as better. This allows us to interpret quality differences expressed in JOD units.
@@ 3.8. JOD regression
(16) $Q_{\textrm{JOD}}=10-\alpha_{\textrm{JOD}}\,(D_{\textrm{pooled}})^{\beta_{\textrm{JOD}}}\,,$
@@ 3.9. Image quality
(17) $D_{\textrm{pooled}}^{\textrm{image}}=k_{\text{I}}\left\|w_{c}\left\|\frac{1}{N_{b}^{\nicefrac{{1}}{{\beta_{x}}}}}\left\|\hat{D}_{b,c,f}(\bm{x})\right\|_{\beta_{x},\bm{x}}\right\|_{\beta_{b},b}\right\|_{\beta_{c},c}\,.$
@@
3.10. Visualization — heatmaps and distograms
Figure 11. Examples of distogram visualization of artifacts. The values on the y-axis denote spatial frequency in cpd. The higher (more yellowish) values denote stronger distortion. The low-frequency waveguide nonuniformity (left) is shown to be visible mostly in the base band (BB) and low frequency bands of both achromatic (A-sust) and chromatic (RG, YV) channels. The artifacts shows also as low-frequency flicker in the achromatic transient channel (A-trans). The (per-pixel) light source nonuniformity causes high-frequency shimmer, which is visible as high frequency distortions in the distogram on the right.
@@ 3.10. Visualization — heatmaps and distograms
(18) $H_{b}(\bm{x})=\left\|w_{c}\hat{D}_{b,c,f}(\bm{x})\right\|_{\beta_{c},c}$
@@
Another visualization, which we name a distogram, lets us present video distortions within each channel and spatial frequency band, all in a single diagram. Two examples for two distortions from the XR-DAVID dataset are shown in Fig. 11. The examples show that depending on the characteristics of the distortion, the artifacts will show in either low- or high-frequency bands, in achromatic or chromatic channels. This visualization decomposes the distortions into their frequency components, visual channels, and their time series, which helps interpret and explain how those distortions contributed to the final JOD score.
3.11. Implementation details
@@
4. XR Display Artifact Video Dataset (XR-DAVID)
@@
To address this gap in quantitative data for XR distortions, we enumerated the most relevant artifacts for our study. We placed special emphasis on distortions that have a color or temporal component, as these are not well served by existing metrics and are the focus of ColorVideoVDP. We used this dataset of distortions to conduct a large-scale subjective study, collecting JOD quality scores in a new XR-DAVID video quality dataset [^4] The XR-DAVID dataset can be downloaded from https://doi.org/10.17863/CAM.108210, which is suitable for calibration of ColorVideoVDP.
Figure 12. Physical hardware setup for the XR-DAVID data collection. Participants used a chinrest to fix their distance from the display at 0.73 m (two display heights) and used a 3-key keyboard to record their choices.
4.1. Experimental setup
@@
We selected an Eizo CG3146 professional reference monitor [^5] Eizo CG3146, see https://www.eizo.com/products/coloredge/cg3146/ for detail as our test vehicle. This 31.1” diagonal display has a resolution of $4096\times 2160$, and a contrast of 1,000,000:1 claimed by the manufacturer. The monitor was set to a maximum luminance of 300 cd/m2, with sRGB EOTF, P3 color primaries, and a 60 Hz refresh rate. This display has a built-in colorimeter, which allowed for high-precision daily calibration to ensure accuracy as the study progressed. The distance of the observer was controlled using a chin rest, which was placed so that the effective resolution of the display was 77 pixels per visual degree (see Fig. 12).
@@
We conducted the study using paid external participants. After two pilots (N=5 each), 77 naïve users took part in 1 hour long sessions of our main study, which included training, data gathering, and a short break half-way through. All participants signed informed consent forms, and the study was approved by an external ethics committee. Participants were screened via an Ishihara color test (Hardy et al., 1945) to ensure normal color vision.
@@
Participants performed a side-by-side pairwise comparison task answering which of the two versions of a video was less distorted. This method was preferred to alternatives (e.g., direct ratings on the mean-opinion-score scale) as pairwise comparisons have been shown to produce more accurate results (Zerman et al., 2018) by simplifying the task in each trial. Active sampling using the ASAP method (Mikhailiuk et al., 2021) was used to optimize information gain from each comparison, making it possible to explore a large number of artifacts without prohibitively increasing the number of pairwise comparisons. We allowed comparisons across different distortion types and levels but not across different video contents. A training session preceded the main experiment. In that session, the participants compared each reference video to its distorted version. A different distortion type was used in each trial. The training session was meant to familiarize the participants with the reference content and distortions. The results of the main experiment were scaled to a unified perceptual scale in just-objectionable-difference units (JODs) using the pwcmp software suite (Perez-Ortiz et al., 2019).
4.2. Distortions
@@
Figure 13. MTF Curves corresponding to the Blur artifact at level 1, 2, and 3.
@@
AR displays with diffractive waveguides such as Microsoft Hololens, Magic Leap, and Snap Spectacles (Ooi and Dingliana, 2022) exhibit a characteristic spatially-varying color nonuniformity. This nonuniformity is typically low-frequency (less than  1 cycle per degree). In addition, it is heavily dependent on pupil position within the eye box, which can vary depending on the user’s fixation and eye movements. In particular, if a user keeps their gaze fixed on an object in the AR content while moving their head (engaged in what is typically termed a vestibulo-ocular reflex or VOR movement), they may notice its color shifting as the position of the eye changes in relation to the waveguide. In order to simulate the artifact, first, two nonuniformity patterns were obtained from empirical waveguide transmission data. Second, base variation maps were generated for the two patterns by cleaning noise in data and normalizing variations across channels. The variations were also scaled such that they appear to have similar magnitude across the two patterns. Finally, the amplitude of variations in the base map was varied to produce different artifact levels. The multipliers of the amplitude variations were 0.15, 0.3, and 0.4 for artifact levels 1, 2, and 3 respectively. To simulate changes with eye position, the nonuniformity pattern transitions from the first base pattern to the second and back over the course of the video. This transition is repeated 4 times for level 2, and 7 times for level 3. Fig. 14 illustrates this temporally varying waveguide nonuniformity artifact.
Figure 14. Examples of waveguide nonuniformity level 3 at different frames, using Business scene (cropped for better visibility). nonuniformity changes from one pattern (left) to another (right).
@@
In AR displays suffering from waveguide nonuniformity across different exit pupil positions, as described above, a dynamic correction algorithm can be implemented to invert the distortion, reducing spatial nonuniformity. This algorithm would be dependent on gaze location, thus relying on estimated positions from eye trackers, which may be inaccurate. Namely, inaccurate eye tracker readings will cause a wrong estimate of the pupil position and temporal color artifacts (by either not correcting perfectly for distortions or performing an imprecise correction for a distortion pattern that is not present). To simulate this, the precision error was randomly generated per frame from a Gaussian distribution with a standard deviation of 0.27°, 0.39°, and 0.58° for artifact levels 1, 2, and 3 respectively. Simulated precision errors are in alignment with existing VR headsets currently in the market (Schuetz and Fiehler, 2022).
@@ Results
Figure 15. The results of our display artifact study aggregated over all users and scenes are shown. Quality scores are expressed in JODs, counting down from 10 by convention. Values above the nominal reference are the result of chance selection for artifacts that appear very subtle. Error bars represent 95% confidence intervals, obtained via bootstrapping.
5. Metric validation
Table 2. Datasets used for training and testing.
@@ 5. Metric validation
UPIQ (Mikhailiuk et al., 2022) Train & test SDR/HDR images 84 4159 34 distortion types
LIVE HDR (Shang et al., 2022) Test HDR video 21 210 H.265, bicubic upscaling
LIVE VQA (Seshadrinathan et al., 2010) Test HDR video 10 150 H.264, MPEG-2, transmission
KADID-10k (Lin et al., 2019) Test SDR images 81 10125 25 distortion types
@@
We split the datasets into those used solely for testing and those used for both training and testing. The five selected datasets are listed in Table 2. The two datasets used for both training and testing are XR-DAVID, explained in detail in Sec. 4, and UPIQ (Mikhailiuk et al., 2022). We selected the UPIQ dataset because it contains a large collection of both SDR and HDR images (over 4000) and is scaled in JOD units, similar to XR-DAVID. To test our metric on unseen datasets (cross-dataset validation), we chose LIVE HDR because it is a modern dataset representative of video streaming applications, KADID-10k because of its size (over 10k images), and LIVE VQA because it is widely used to test video quality metrics. Because KADID-10k and a part of UPIQ were collected in uncontrolled crowdsourcing experiments, we had to approximate the display specification and viewing conditions for those. We avoided video datasets that were collected in uncontrolled conditions (e.g., KonViD-1k).
@@
Training a video quality metric is problematic as a very large amount of data is used to predict a single quality value. For example, 60 frames of 4K video requires the processing of almost 500 million pixels to infer just a single quality score. This makes gradient computation (through backpropagation) infeasible because of the memory requirements. Previous work dealt with this problem in several ways: some metrics were designed to extract low dimensional features from a video and then train a regression mapping those features to quality scores (Li et al., 2016b). This approach, however, does not allow training or tuning the feature extraction stage. Another group of methods operated on patches of limited resolution (e.g. 64$\times$64), assuming that the quality is the same across the entire image (Prashnani et al., 2018). This assumption, however, can be easily proved wrong for localized distortions or for content in which the effect of contrast masking varies across an image. Other metrics used numerical gradient computation (Mantiuk et al., 2021), which, however, becomes too expensive when optimizing a large number of parameters.
@@
For the end-to-end training, we computed the full analytical gradient of the remaining parameters using two techniques. Firstly, we utilized gradient checkpointing (Chen et al., 2016) in PyTorch, by rerunning a forward pass for each checkpointed segment during backward propagation. This technique trades some speed for reduced memory requirements. We inserted a checkpoint after each block of frames, where the block size was determined based on the available GPU memory. Secondly, during training, we randomly sampled 0.5-second-long sequences from each video clip (full sequences were used for testing). This approach improved training time and introduced a form of data augmentation. The feature-based and end-to-end training were run in an interleaved manner, with 50 epochs of (fast) feature-based training followed by a single epoch of end-to-end training. This lets us jointly train all the parameters of the metric. The trained parameters of the metric can be found in Table 3.
Table 3. The trained parameters of the ColorVideoVDP. The parameters of the cross-channel masking can be found in Fig. 9.
Model component Parameters
Contrast sensitivity $S_{\text{corr}}=0.9683$
Masking $p=2.264$, $q_{c}=[1.303,2.889,3.681,3.589]$, $k_{\text{C}}=366.6$, $\sigma_{\text{sp}}=3$,
@@
Pooling $k_{\text{I}}=0.5779$, $w_{c}=[1,1,1,0.8081]$
JOD regression $\alpha_{\text{JOD}}=0.04396$, $\beta_{\text{JOD}}=0.9302$
5.1. Comparison with other metrics
@@
As our datasets include both SDR and HDR content, we need to ensure that those are handled correctly by all the metrics. For metrics that do not work with colorimetric data and instead operate on display-encoded (SDR) pixel values, we ran the evaluations using the original SDR pixel values. In the case of HDR content, we employed the perceptually-uniform transform (PU21) (Mantiuk and Azimi, 2021) to encode the pixel values. The metrics that operate on colorimetric data (FovVideoVDP, HDR-VDP-3) were supplied the absolute luminance values, computed by the display model (Sec. 3.1). The metrics that account for the viewing distance or the resolution in pixels-per-degree (e.g., HDR-VDP-3, HDR-FLIP) were supplied with the correct values.
@@
The results, shown in Fig. 16, indicate a substantial gain in performance of ColorVideoVDP over the second-best metric — VMAF. Image metrics that consider color — FSIMc and VSI — performed better than expected on XR-DAVID, but rather poorly on LIVEHDR video dataset. The color difference formulas and their spatial extensions performed worse than image and quality metrics. The non-reference metrics — NIQE, PIQE, and BRISQUE — show a very small correlation with subjective judgments. Other performance indices (SROCC and PLCC) and detailed results can be found in the HTML reports included in the supplementary. Overall, while it is possible to find a metric that performs well on a selected dataset, such as STRRED on LIVEVQA (LIVEVQA was used to calibrate and test STRRED), ColorVideoVDP offers good performance across all datasets with a wide variety of distortions and content. This could be explained by no other metric offering the same set of capabilities as ColorVideoVDP; only our metric models spatiotemporal color vision and accounts for the display model.
Figure 16. ColorVideoVDP compared with existing metrics, listed in Table 1. The prediction errors are reported for testing datasets (LIVEVQA, LIVEHDR and KADID) and the test portion of UPIQ and XR-DAVID. The error bars denote 95% confidence intervals. Other performance metrics (SROCC and PLCC) and detailed reports can be found in the supplementary HTML report.
5.2. Ablations
Figure 17. Ablation studies on the contrast encoding and masking models used in ColorVideoVDP. The notation is the same as in Fig. 16. The masking models used in the ablations are explained in the supplementary.
We grouped ablation studies into those used to determine suitable contrast encoding and masking models, and into those used to test the importance of each component of the metric. The metric parameters were refitted for each ablation, as explained above. First, we tested a combination of two contrast encodings and three masking functions, all explained in detail in the supplementary document. The results of those ablations, shown in Fig. 17, clearly indicate that the mutual masking model with multiplicative contrast encoding offers much better performance than the alternatives. The results indicate that the multiplicative contrast encoding is necessary for mutual masking and transducer models. As mentioned in Sec. 3.6, such encoding can unify results across different spatial frequencies (Daly, 1993) and color directions (Cass et al., 2009; Switkes et al., 1988). Although the contrast transducer is one of the best-established masking models (Watson and Solomon, 1997), which accounts for facilitation and performs well on selected datasets (Alam et al., 2014), it did not perform well as a part of the quality metric. The similarity formula exhibits masking properties and is used in many metrics, such as (MS-)SSIM (Wang et al., 2003), but it did not result in acceptable performance when integrated into our metric.